# Open-vocabulary Temporal Action Localization using VLMs

### v3

@@ V Discussion and Limitation
In this paper, we proposed a method for open-vocabulary video action localization using VLMs. Despite its straightforward pipeline, benchmark evaluations demonstrated reasonable performance, both for third-person and first-person videos. The effectiveness of this approach diminishes with increased video step lengths, and it is outperformed by established learning-based methods. Nevertheless, this method offers significant advantages: it eliminates the need for data collection or training and can extract actions specified by open-vocabulary free-text queries, thereby enhancing its adaptability to diverse applications such as video annotation and video editing.
The comparison of different prompting styles has revealed several limitations of this method. First, the optimal number of frames for sampling depends on a trade-off between spatial and temporal resolution in action recognition. This number may vary based on the video’s length and the nature of the actions, potentially requiring adjustments for each use case. Second, the method’s performance is negatively impacted by the length of action sequences and by numerical annotations that interfere with image clarity. These issues highlight opportunities to improve the method by adjusting the placement of number renderings or by modifying the video length beforehand.
@@ VI Conclusion
This paper introduces a novel approach for open-vocabulary video action localization leveraging the PIVOT methodology, which eliminates the need for data collection and model training. Testing with benchmark datasets has demonstrated that the proposed method offers promising capabilities in action localization without the constraints of predefined vocabularies. Considering several limitations of this method, future work will focus on refining visual prompting techniques and optimizing the system to enhance its applicability across a broader range of video analysis tasks, from automated video labeling to advanced robotics training.

### v4

@@ V Discussion and Limitation
In this paper, we propose a method for VLM-empowered OVAL, using GPT-4o as an example of VLMs. Despite its straightforward iterative pipeline, our method demonstrated reasonable performance in both third-person and first-person videos. The results showed reasonable performance in task-transition estimation, with limitations such as decreased performance on videos with longer action sequences. Furthermore, the method demonstrated performance comparable to state-of-the-art zero-shot action localization. Importantly, the unique advantage of the proposed method lies in its training-free nature, eliminating the need for data collection or training. This allows for immediate and effortless application to new vocabularies and dataset domains, offering adaptability to diverse applications such as video annotation and editing.
The comparison of different prompting styles has revealed several limitations of this method. First, the optimal number of frames for sampling depends on a trade-off between spatial and temporal resolution in action recognition. This number may vary based on the video’s length and the nature of the actions, potentially requiring adjustments for each use case. Second, the method’s performance is negatively impacted by the length of action sequences and by numerical annotations that interfere with image clarity. These issues highlight opportunities to improve the method by adjusting the placement of number renderings or by modifying the video length beforehand, as employed in the zero-shot experiment. Lastly, since our method requires multiple queries to the VLM, it takes a longer time to recognize actions compared to typical learning-based methods. This issue could potentially be mitigated by using a more lightweight or local model that does not rely on network communication.
@@ VI Conclusion
This paper introduces a novel approach for open-vocabulary video action localization leveraging the PIVOT methodology, which eliminates the need for data collection and model training. Testing with benchmark datasets has demonstrated that the proposed method offers promising zero-shot capabilities without the need for model pre-training. Considering several limitations of this method, future work will focus on refining visual prompting techniques and optimizing the system to enhance its applicability across a broader range of video analysis tasks, including automated video labeling and advanced robotics training.